OpenStudy (erinweeks):

What is the volume of 60.0 g of ether if the density of ether is 0.70 g/mL? A. 1.2 × 10–2 mL B. 86 mL C. 2.4 × 10–2 mL D. 42 mL.

OpenStudy (jamesj):

Write p for density, m for mass and V for volume. Then by definition of density p = m/V Therefore V = m/p You are given m and you given p. Now calculate V.

OpenStudy (erinweeks):

im confused could you show me the steps

OpenStudy (erinweeks):

i got 85.714 so would it be B?

OpenStudy (jamesj):

Looks about right.

OpenStudy (jtfan2000):

42